Given the following budget line, (M = PfF + PsS) show the following by reformulating the equation. Assume F (food) is on the horizontal axis and S (shelter) is on the vertical axis.

A. The vertical intercept is ________.
B. The horizontal intercept is ________.
C. The slope of the budget line is ________.
Answer #1

A.The vertical intercept=M/Ps

B.Horizontal intercept=M/Pf

C.Slope of the budget line=-Pf/Ps

The intercept is determined by the expenditure of the consumer when exhausting all income on the purchase of the good on the respective axis.